Hi. I'm trying to help a friends little sister with her homework but is been so long since Ive done this stuff I want to check my answer to make sure its right. the question is f + 3 11/12=18. Anyone who can help, it would be greatly appreciated.
Answer by josmiceli(19441)

f + 3 11/12 = 18
convert both sides to fractions
f + 47/12 = (18 * 12)/12
f = 216/12 - 47/12
f = 169/12